OpenStudy (anonymous):

How many grams of HCl are produced when 2.5 kg C12H4Cl6 react?

OpenStudy (shikamaru11):

@shelby.lane ma'm with whom it reacts or just decomposes???

OpenStudy (shikamaru11):

by balancing the equation..... 2C12-H4-Cl6 + 23O2 + 2H2O = 24CO2 + 12HCl n(PCB) = m(PCB) / MW(PCB) = 2500 / 361 = 6.93 mol According to equation n(HCl) = 6n(PCB) = 6 x 6.93 = 41.58 mol m(HCl) = n(HCl) x MW(HCl) = 41.68 x 36.5 = 1517.7 g Answer: m(HCl) = 1517.7 g = 1.5 kg

OpenStudy (shikamaru11):

@shelby.lane

OpenStudy (anonymous):

Yes! SO why isn't it taking my answer?

OpenStudy (shikamaru11):

@shelby.lane Don't know but it's rite one as you see